OpenStudy (anonymous):

Please help ASAP! What is the solution to y = 2x + 3 for the replacement set x = {1.3}?

OpenStudy (anonymous):

@bohotness @KyanTheDoodle @King.Void. @purple_pink

OpenStudy (purple_pink):

hi

OpenStudy (anonymous):

hi, please help!

OpenStudy (purple_pink):

Substitute x = 1.3 into the equation " y = 2x+3" So basically where you see "x" in y = 2x + 3, replace the x with 1.3 y = 2 (1.3) + 3 y = ? Hope this helps!!:)

OpenStudy (anonymous):

...i dont know the answer from that exactly :P im dumb with math

OpenStudy (purple_pink):

Simple. Substitue "1.3" for x in the given equation y = 2x + 3. 2x + 3 = y 2 * 1.3 + 3 = y 2.6 + 3 = y 5.6 = y The solution to the equation with the replacement set {1.3} for x is {5.6} = y.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

oh wait @purple_pink i need two parts

OpenStudy (purple_pink):

ok

OpenStudy (anonymous):

What is the solution to y = 2x + 3 for the replacement set x = {1.3}? (__ , ___)

OpenStudy (purple_pink):

uh hld on

OpenStudy (anonymous):

@purple_pink

OpenStudy (anonymous):

k

OpenStudy (purple_pink):

i tagged you